# Early-Exit Graph Neural Networks

### v1

@@ B.1 The Gumbel Distribution and the Gumbel-Softmax Temperature
Here, $\nu_{0}$ sets the minimum inverse temperature, thereby controlling the upper bound of the temperature. In our setup, we implement $f_{\nu}$ as in Equation 14, with $\text{g}=\text{GNN}(\cdot)$ for node classification (see Algorithm 1) and $\text{g}=\text{MLP}(\cdot)$ for graph classification (see Algorithm 2). The confidence estimator $f_{c}$ is defined as $\text{g}(\mathbf{H}^{l})$, choosing g based on the task, similarly to $f_{\nu}$.
@@ Appendix C Proofs of Section 3
The following exposition share similarities with Chang et al. [2019], Gravina et al. [2023], Di Giovanni et al. [2023].
@@
where we apply the spectral decomposition of $J$, and get its eigenvectors $\mathbf{T}$, and eigenvalues $\mathbf{\Lambda}=diag(\lambda_{i})$. As denoted in Gravina et al. [2023], to guarantee the stability of the system, we require the $Re(\lambda_{i})\leq 0,\forall i=1,\dots,d$. However, when the eigenvalues are less than zero, we can lose the information of the previous node representations, leading to a dissipative behavior. For this reason we want $Re(\lambda_{i})=0,\forall i=1,\dots,d$, having only imaginary eigenvalues. In the case of Equation (7), the Jacobian has only imaginary eigenvalues and takes the form of
@@
Here, the choice of the antisymmetric multiplication by $\mathbf{\Omega}-\mathbf{\Omega}^{\top}$ guarantees the presence of imaginary eigenvalues, and left-multiplying it by a diagonal matrix does not change the nature of the overall spectrum. This was proved in Gravina et al. [2023]. Our Jacobian instead becomes:
@@
Since we assume that the derivative of $\sigma_{1}$ is bounded, the assumption that $J^{t}$ stays constant can be valid. Hereafter, we show this proof, similarly to Gravina et al. [2023].
@@
$\mathbf{A}$ is a diagonal matrix with non-negative entries, let us denote by $\mathcal{I}\in[d]$ the set of indices s.t. =$\mathcal{I}=\{i\in[d]:\mathbf{A}_{ii}=0\}$. If $\mathbf{v}$ eigenvector of $\mathbf{AB}$ then $\mathbf{v}_{j}=0\;\forall j\in\mathcal{I}$. Indeed, we need for each entry $k$, $\sum_{i}\mathbf{AB}_{ki}v_{i}=\lambda v_{k}$; for $j$ in $\mathcal{I}$ the row the $j-$th row of $\mathbf{AB}$ is 0. So we have $\lambda v_{j}=0$, since we assumed $\lambda\neq 0,$ this implies $v_{j}=0$.
@@
We have already observed that if $\mathbf{v}$ eigenvector of $\mathbf{AB}$, it has entries in $\mathcal{I}$ equal to 00. If $\mathbf{AB}\mathbf{v}=\lambda\mathbf{v}$, using (26),
@@
Now, what we aim to understand is whether Equation 15 can also induce attraction and repulsion edge-wise as in Di Giovanni et al. [2023]. Our assumptions are the same as the previous proof, but we also assume $\sigma_{2}$ to be a point-wise activation function that returns only positive values, which do not violate the assumption of the previous proof.
@@
$=-\mathbf{Z}^{t}\sigma(\mathbf{Z}^{t}+\mathbf{Y}^{t})$ (41)
@@
$\mathbf{Z}^{t}_{i}\cdot\sigma(\mathbf{Z}^{t}_{i}+\mathbf{Y}^{t}_{i})\geq 0\hskip 14.22636pt\forall i=1,...nm^{\prime},$ (42)
@@
ReLU, or $tanh$, does not satisfy this condition, thus, we elaborate on this step by imposing the following constraints.
$\frac{dE_{\theta}(\mathbf{H}^{t})}{dt}=-\mathbf{Z}^{t}\sigma(\mathbf{Z}^{t}-\text{ReLU}(\mathbf{Y}^{t}))$ (43)
@@
The conditions in Equation (42) are always satisfied, since when $\mathbf{Z}^{t}_{i}<0$, we have that everything is 00. Otherwise, we consider two cases, namely when $\mathbf{Z}^{t}_{i}>\text{ReLU}(\mathbf{Y}^{t}_{i})$ and $\mathbf{Z}^{t}_{i}<\text{ReLU}(\mathbf{Y}^{t}_{i})$. In the former, we have that $\text{ReLU}(\text{tanh}(\cdot))$ is always positive, and also $\mathbf{Z}^{t}_{i}$. In the latter case, we have a negative argument, and the Equation goes to 00. Generally, we could satisfy these conditions for any choice of $\sigma_{1}(x)\geq 0,\forall x\in\mathbb{R}$, and also for $\sigma_{2}\geq 0,\forall x\in\mathbb{R}$. We conclude that through our assumptions, Equation (15) can induce attraction and repulsion edge-wise.
@@
Proof of the theorem. Theorem 3.4, is simple to demonstrate, since $-\sigma_{2}(\mathbf{B}\mathbf{E}\mathbf{W}_{e})$ is not dependent on the node features and can be encompassed in the $\phi_{W_{s}}(\mathbf{H}^{t},\Gamma(i))$ term of Equation (16), and repeat the same proof. According to this, Equation (4) is stable and non-dissipative. To prove that it induces attraction and repulsion edge-wise, we can take Equation (39), and now specify that $\mathbf{Y}^{t}=-\sigma_{2}((\mathbf{\Omega}_{as}^{\top}\otimes\mathbf{I}_{n})\text{vec}(\mathbf{H}^{t}))-\sigma_{2}((\mathbf{E}^{\top}\otimes\mathbf{I}_{n})\text{vec}(\mathbf{B}))$. We will derive the same result, since each element of $\mathbf{Y}^{t}$ is less than or equal to 0.
@@ Appendix D Early-Exit Graph Neural Networks Algorithm for Graph Classification
In Algorithm 1, we describe how the Straight-Through Gumbel-Softmax estimator is integrated with SAS-GNN to allow each node to decide, at every layer, whether to exit or continue processing. However, since we also address graph classification, we refine this procedure to enable early-exit decisions at the graph level. Specifically, instead of applying the Gumbel-Softmax to individual nodes as in Algorithm 1, we apply it to the entire graph—so that the model determines whether a graph’s representation is sufficiently complete to make a prediction. This modification ensures that exit decisions align with the prediction target: nodes for node classification, graphs for graph classification. To maintain consistency, we define the "agent"—the entity responsible for triggering the exit—as the unit on which the prediction task is performed: nodes for node-level tasks and graphs for graph-level tasks. While it is theoretically possible to perform graph classification with node-level exits (e.g., by aggregating node predictions), we leave such extensions for future work. The refined algorithm follows.
@@
1:Initialize $\mathcal{G}=(\mathbf{H}^{0},\mathbf{\bar{A}}),\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d},L,f_{c},f_{\nu},f_{e},\text{Pool},\mathbf{W}_{s},\mathbf{\Omega}_{as},\sigma_{1},\sigma_{2},\nu_{0}$
2:for $l=0$ to $L$ do
3:     $C^{l}\leftarrow f_{c}(\text{Pool}(\mathbf{H}^{l}))$
4:     $\nu^{l}\leftarrow f_{\nu}(\text{Pool}(\mathbf{H}^{l},\nu_{0}))$
5:     $\mathbf{c}^{l}\leftarrow gumbel\_softmax(C^{l},\nu^{l})$
6:     $\tau^{l}\leftarrow\mathbf{c}^{l}(\texttt{0})$
7:     $\mathbf{H}^{l+1}\leftarrow\mathbf{H}^{l}+\tau^{l}\sigma_{1}(-\sigma_{2}(\mathbf{H}^{l}\mathbf{\Omega}_{as})+f_{e}(\mathbf{E})+\mathbf{\bar{A}}\mathbf{H}^{l}\mathbf{W}_{s})$
8:     if $\arg\max\{\mathbf{c}^{l}\}=\texttt{1}$ then
9:         $\mathbf{Z}\leftarrow pool(\mathbf{H}^{l})$
10:         return $\mathbf{Z}$
11:$\mathbf{Z}\leftarrow pool(\mathbf{H}^{L})$
12:return $\mathbf{Z}$
@@ E.1 Datasets
Datasets $N$ $|\mathcal{E}|$ $d$ $|C|$ $\xi_{edge}$ $\xi_{adj}$
@@
Table 5: Dataset Information
The dataset released by [Dwivedi et al., 2023] instead has the statistics reported in Tables 6 and 7.
@@ E.2 Baselines
Baselines for heterophilic graphs. We consider the set of heterophilic graphs introduced in Platonov et al. [2023]. We test our models in heterophilic settings, since these are scenarios that typically cause over-smoothing, and we want to gauge in practice how they compare against existing models. We briefly describe the baselines used for comparison against SAS-GNN and EEGNN in node classification. From the benchmark introduced in [Platonov et al., 2023], we include the reported performances of GCN, GraphSAGE [Kipf and Welling, 2017, Hamilton et al., 2017], GAT [Veličković et al., 2018], and GT [Dwivedi and Bresson, 2021a]. Additionally, we consider GAT-sep and GT-sep, two variants of these models specifically designed for heterophily. Even though these baselines seem outdated, it is known that the experiments in [Platonov et al., 2023] revealed that classic models outperform methods specialized for heterophily. More recently, Finkelshtein et al. [2024] introduced the Cooperative Graph Neural Networks (Co-GNNs) model, which we also include in our evaluation. This family of models is also more expressive than classic message-passing neural networks [Xu et al., 2019], thanks to an asynchronous message-passing.
@@
Baselines for Long Range Graph Benchmark. We compare SAS-GNN and EEGNN on 3 datasets from LRGB Dwivedi et al. [2023]. These datasets were collected to assess how much GNNs can preserve information exchanged among distant nodes. We take advantage of these datasets to test how much our model performs compared to other methods. Among the baselines, we compare Classic MPNNs: GCN Kipf and Welling [2017], GIN Xu et al. [2019], GINE Hu et al. [2020], GatedGCN Bresson and Laurent [2018], which is the category that encompasses SAS-GNN and EEGNN. These methods need to process $l$ message-passing steps to allow nodes distant $l$ to communicate. Graph Transformers: Dwivedi and Bresson [2021a], SAN Kreuzer et al. [2021], and GraphGPS Rampášek et al. [2023], this category uses self-attention to all the nodes, so it allows nodes to reach every part of the graph, but it has a quadratic complexity w.r.t $|\mathcal{V}|$. Rewiring methods: DIGL Gasteiger et al. [2022], MixHop Abu-El-Haija et al. [2019], and DRew Gutteridge et al. [2023], these approaches modify the graph topology during the message-passing, in this way messages can be delivered through paths that did not exist before, and so long-range interactions are fostered. Asynchronous Message-Passing methods: Co-GNN Finkelshtein et al. [2024], AMP Errica et al. [2023]. These can be seen similarly to graph rewiring, since the computational graph is different from the input graph, as well as rewiring methods. However, here there exists the interpretation of nodes that decides what messages to filter, rather than rewiring the graph.
@@
We test our approach only on 3 out of 5 datasets, because of computational constraints. These are Peptides-func, Peptides-struct, and Pascal-VOC.

### v2

@@ B.1 The Gumbel Distribution and the Gumbel-Softmax Temperature
Here, $\nu_{0}$ sets the minimum inverse temperature, thereby controlling the upper bound of the temperature. In our setup, we implement $f_{\nu}$ as in Equation 18, with $\text{g}=\text{GNN}(\cdot)$ for node classification (see Algorithm 1) and $\text{g}=\text{MLP}(\cdot)$ for graph classification (see Algorithm 2). The confidence estimator $f_{c}$ is defined as $\text{g}(\mathbf{H}^{l})$, choosing g based on the task, similarly to $f_{\nu}$. We implemented $f_{\nu}$, and $f_{c}$ using SAS-GNN in the ECHO benchmark experiments, while we used MeanGNN [Finkelshtein et al., 2024] otherwise.
@@ Appendix C Proofs of Section 3
The following exposition share similarities with [Chang et al., 2019, Gravina et al., 2023, Di Giovanni et al., 2023].
@@
where we apply the spectral decomposition of $J$, and get its eigenvectors $\mathbf{T}$, and eigenvalues $\mathbf{\Lambda}=diag(\lambda_{i})$. As denoted in [Gravina et al., 2023], to guarantee the stability of the system, we require the $Re(\lambda_{i})\leq 0,\forall i=1,\dots,d$. However, when the eigenvalues are less than zero, we can lose the information of the previous node representations, leading to a dissipative behavior. For this reason we want $Re(\lambda_{i})=0,\forall i=1,\dots,d$, having only imaginary eigenvalues. In the case of Equation (9), the Jacobian has only imaginary eigenvalues and takes the form of
@@
Here, the choice of the antisymmetric multiplication by $\mathbf{\Omega}-\mathbf{\Omega}^{\top}$ guarantees the presence of imaginary eigenvalues, and left-multiplying it by a diagonal matrix does not change the nature of the overall spectrum. This was proved in [Gravina et al., 2023]. Our Jacobian instead becomes:
@@
Since we assume that the derivative of $\sigma_{1}$ is bounded, the assumption that $J^{t}$ stays constant can be valid. Hereafter, we show this proof, similarly to [Gravina et al., 2023].
@@
$\mathbf{A}$ is a diagonal matrix with non-negative entries, let us denote by $\mathcal{I}\in[d]$ the set of indices s.t. $\mathcal{I}=\{i\in[d]:\mathbf{A}_{ii}=0\}$. If $\mathbf{v}$ eigenvector of $\mathbf{AB}$ then $\mathbf{v}_{j}=0\;\forall j\in\mathcal{I}$. Indeed, we need for each entry $k$, $\sum_{i}\mathbf{AB}_{ki}v_{i}=\lambda v_{k}$; for $j$ in $\mathcal{I}$ the row the $j$-th row of $\mathbf{AB}$ is 0. So we have $\lambda v_{j}=0$, since we assumed $\lambda\neq 0,$ this implies $v_{j}=0$.
@@
We have already observed that if $\mathbf{v}$ eigenvector of $\mathbf{AB}$, it has entries in $\mathcal{I}$ equal to 0. If $\mathbf{AB}\mathbf{v}=\lambda\mathbf{v}$, using (30),
@@
Now, what we aim to understand is whether Equation 19 can also induce attraction and repulsion edge-wise as in [Di Giovanni et al., 2023]. Our assumptions are the same as the previous proof, but we also assume $\sigma_{2}$ to be a point-wise activation function that returns only positive values, which do not violate the assumption of the previous proof.
@@
$=-\mathbf{Z}^{t}\sigma_{1}(\mathbf{Z}^{t}+\mathbf{Y}^{t})$ (45)
@@
$\mathbf{Z}^{t}_{i}\cdot\sigma_{1}(\mathbf{Z}^{t}_{i}+\mathbf{Y}^{t}_{i})\geq 0\hskip 14.22636pt\forall i=1,...nm^{\prime},$ (46)
@@
ReLU, or tanhtanh, does not satisfy this condition, thus, we elaborate on this step by imposing the following constraints.
$\frac{dE_{\theta}(\mathbf{H}^{t})}{dt}=-\mathbf{Z}^{t}\sigma_{1}(\mathbf{Z}^{t}-\text{ReLU}(\mathbf{Y}^{t}))$ (47)
@@
The conditions in Equation (46) are always satisfied, since when $\mathbf{Z}^{t}_{i}<0$, we have that everything is 0. Otherwise, we consider two cases, namely when $\mathbf{Z}^{t}_{i}>\text{ReLU}(\mathbf{Y}^{t}_{i})$ and $\mathbf{Z}^{t}_{i}<\text{ReLU}(\mathbf{Y}^{t}_{i})$. In the former, we have that $\text{ReLU}(\text{tanh}(\cdot))$ is always positive, and also $\mathbf{Z}^{t}_{i}$. In the latter case, we have a negative argument, and the Equation goes to 0. Generally, we could satisfy these conditions for any choice of $\sigma_{1}(x)\geq 0,\forall x\in\mathbb{R}$, and also for $\sigma_{2}\geq 0,\forall x\in\mathbb{R}$. We conclude that through our assumptions, Equation (19) can induce attraction and repulsion edge-wise.
@@
Proof of the theorem. Theorem 3.4, is simple to demonstrate, since $\sigma_{2}(\mathbf{B}\mathbf{E}\mathbf{W}_{e})$ is not dependent on the node features and can be encompassed in the $\phi_{W_{s}}(\mathbf{H}^{t},\Gamma(i))$ term of Equation (20), and repeat the same proof. According to this, Equation (4) is stable and non-dissipative. To prove that it induces attraction and repulsion edge-wise, we can take Equation (43), and now specify that $\mathbf{Y}^{t}=-\sigma_{2}((\mathbf{\Omega}_{as}^{\top}\otimes\mathbf{I}_{n})\text{vec}(\mathbf{H}^{t}))-\sigma_{2}((\mathbf{E}^{\top}\otimes\mathbf{I}_{n})\text{vec}(\mathbf{B}))$. We will derive the same result, since each element of $\mathbf{Y}^{t}$ is less than or equal to 0.
@@ Appendix D Early-Exit Graph Neural Networks Algorithm for Graph Classification
Neural Adaptive Step for Graph Classification. In Algorithm 1, we describe how the Straight-Through Gumbel-Softmax estimator is integrated with SAS-GNN to allow each node to decide, at every layer, whether to exit or continue processing. However, since we also address graph classification, we refine this procedure to enable early-exit decisions at the graph level. Specifically, instead of applying the Gumbel-Softmax to individual nodes as in Algorithm 1, we apply it to the entire graph, so that the model determines whether a graph’s representation is sufficiently complete to make a prediction. This modification ensures that exit decisions align with the prediction target: nodes for node classification, graphs for graph classification. To maintain consistency, we define the "agent"—the entity responsible for triggering the exit—as the unit on which the prediction task is performed: nodes for node-level tasks and graphs for graph-level tasks. While it is theoretically possible to perform graph classification with node-level exits (e.g., by aggregating node predictions), we leave such extensions for future work. The refined algorithm follows.
@@
[…]
@@
Relation to Adaptive-Step in Runge-Kutta solvers. Our framework is conceptually related to adaptive-step ODE solvers such as DOPRI [Dormand and Prince, 1980b]. However, our approach differs in that a neural network determines the stopping condition, whereas DOPRI solvers rely on integration error tolerance. Although integrating these methods may be beneficial, it is outside the scope of this work.
Mitigation of Underreaching. EEGNN is inherently robust to under-reaching by treating the total depth $L$ as a computational budget rather than a fixed architectural constraint. As demonstrated in Figure 3, increasing $L$ does not degrade performance, as the model naturally exits early when deep processing is unnecessary. Crucially, the Neural Adaptive Step, which is driven exclusively by the task loss, tend to align with the effective depth with the problem radius. For tasks requiring long-range propagation, such as sssp (Figure 2) and ecc (Figure 17), the model learns to delay exits, utilizing the full depth to capture global dependencies. Conversely, on tasks with local signals like Peptides-func and Peptides-struct (Figure 26), the distribution shifts towards shallow layers. Unlike prior works that learn a single fixed depth for the entire dataset [Errica et al., 2023], our approach adapts depth dynamically per sample. It is true that our method is not tailored for mitigating underreaching, but [Errica et al., 2023] is complementary to our effort, as we could use methodologies learn the budget $L$ without changing our model.
@@ E.1 Datasets
Table 5: Dataset Information
Datasets NN $|\mathcal{E}|$ $d$ $|C|$ $\xi_{edge}$ $\xi_{adj}$
@@
Table 6: ECHO benchmarks statistics.
Dataset # Graphs Avg Nodes Avg Deg. Avg Edges Avg Diam # Node Feat # Edge Feat # Tasks
Graph Topologies 10,080 $83.69\pm 66.24$ $2.53\pm 1.19$ $211.63\pm 209.39$ $28.50\pm 6.92$ 2 None 3
line 1,680 $75.60\pm 27.32$ $2.37\pm 0.10$ $90.10\pm 33.89$ $28.50\pm 6.92$ 2 None 3
ladder 1,680 $56.52\pm 13.82$ $2.92\pm 0.02$ $82.54\pm 20.72$ $28.50\pm 6.92$ 2 None 3
grid 1,680 $193.10\pm 93.10$ $2.95\pm 0.12$ $288.32\pm 145.29$ $28.50\pm 6.92$ 2 None 3
tree 1,680 $60.42\pm 17.17$ $1.96\pm 0.01$ $59.42\pm 17.17$ $28.50\pm 6.92$ 2 None 3
caterpillar 1,680 $34.71\pm 7.96$ $1.94\pm 0.02$ $33.71\pm 7.96$ $28.50\pm 6.92$ 2 None 3
lobster 1,680 $81.79\pm 25.46$ $1.97\pm 0.01$ $80.79\pm 25.46$ $28.50\pm 6.92$ 2 None 3
In Table 6, we report the statistics associated with the ECHO benchmark. In particular we describe the input node and edge features, the graph structural properties such as average degree, average edges, and average diameter, and we report the associated graph topology. As remarked in the main paper the topologies are structured in a way that the tasks can be solved only through a correct and effective long-range reasoning. This is due to the nature of the task but also their structure. These graphs contains topological bottlenecks because of the type of generated topology, and this is a good way of measuring how models behave in scenario where topological over-squashing, and also computational over-squashing [Arnaiz-Rodriguez and Errica, 2025]. combined may arise. The tasks that are considered are Single-Source-Shortest Path (sssp), which predicts the shortest path across all the nodes, then we have Eccentricity (ECC), builds on this by computing the longest shortest path from each node. Finally Diameter (diam), is a task where we look for the longest shortest path between any two nodes. All of these tasks require to span information all over the graph, implying global understanding, and long-range reasoning form the modle. This benchmark is ad-hoc for long-range interaction w.r.t. the existing and widely adopted LRGB [Dwivedi et al., 2023].
The dataset released by [Dwivedi et al., 2023] instead has the statistics reported in Tables 7 and 8.
@@ E.2 Baselines
Baselines for heterophilic graphs. We consider the set of heterophilic graphs introduced in [Platonov et al., 2023]. We test our models in heterophilic settings, since these are scenarios that typically cause over-smoothing, and we want to gauge in practice how they compare against existing models. We briefly describe the baselines used for comparison against SAS-GNN and EEGNN in node classification. From the benchmark introduced in [Platonov et al., 2023], we include the reported performances of GCN, GraphSAGE [Kipf and Welling, 2017, Hamilton et al., 2017], GAT [Veličković et al., 2018], and GT [Dwivedi and Bresson, 2021a]. Additionally, we consider GAT-sep and GT-sep, two variants of these models specifically designed for heterophily. Even though these baselines seem outdated, it is known that the experiments in [Platonov et al., 2023] revealed that classic models outperform methods specialized for heterophily. More recently, Finkelshtein et al. [2024] introduced the Cooperative Graph Neural Networks (Co-GNNs) model, which we also include in our evaluation. This family of models is also more expressive than classic message-passing neural networks [Xu et al., 2019], thanks to an asynchronous message-passing.
Baselines for ECHO Benchmark. We categorize baselines into four groups: (i) standard MPNNs (GCN, GIN [Xu et al., 2019], GCNII [Chen et al., 2020]), here we have the message-passing paradigm where each layer dictates how far information propagates; (ii) rewiring methods as DRew [Gutteridge et al., 2023]; (iii) Graph Transformers (GPS [Rampášek et al., 2023], GRIT [Ma et al., 2023]) leveraging global attention; and (iv) related Graph NODEs (GraphCON [Rusch et al., 2022], A-DGN, SWAN [Gravina et al., 2025], PH-DGN [Heilig et al., 2024]), the class to which SAS-GNN and EEGNN belong. Additionally, we include APGCN [Spinelli et al., 2021], a representative early-exit GNN, to contrast our task-driven mechanism against its budget-aware halting. We also employ GRAFF [Di Giovanni et al., 2023] as a symmetric-only ablation, complementing the antisymmetric dynamics of A-DGN.
@@
Baselines for Long Range Graph Benchmark. We compare SAS-GNN and EEGNN on 3 datasets from LRGB [Dwivedi et al., 2023]. These datasets were collected to assess how much GNNs can preserve information exchanged among distant nodes. We take advantage of these datasets to test how much our model performs compared to other methods. Among the baselines, we compare Classic MPNNs: GCN [Kipf and Welling, 2017], GIN [Xu et al., 2019], GINE [Hu et al., 2020], GatedGCN [Bresson and Laurent, 2018], which is the category that encompasses SAS-GNN and EEGNN. These methods need to process $l$ message-passing steps to allow nodes distant $l$ to communicate. Graph Transformers: [Dwivedi and Bresson, 2021a], SAN [Kreuzer et al., 2021], and GraphGPS [Rampášek et al., 2023], this category uses self-attention to all the nodes, so it allows nodes to reach every part of the graph, but it has a quadratic complexity w.r.t $|\mathcal{V}|$. Rewiring methods: DIGL [Gasteiger et al., 2022], MixHop [Abu-El-Haija et al., 2019], and DRew [Gutteridge et al., 2023], these approaches modify the graph topology during the message-passing, in this way messages can be delivered through paths that did not exist before, and so long-range interactions are fostered. Asynchronous Message-Passing methods: Co-GNN [Finkelshtein et al., 2024], AMP [Errica et al., 2023]. These can be seen similarly to graph rewiring, since the computational graph is different from the input graph, as well as rewiring methods. However, here there exists the interpretation of nodes that decides what messages to filter, rather than rewiring the graph.
@@
We test our approach only on 3 out of 5 datasets, because of computational constraints. These are Peptides-func, Peptides-struct, and Pascal-VOC. Results are averaged across 4 different weight initializations.